A survey of a large number of buyers at a suburban Sams Club gave the proportions of buyers falling in the following income-residential categories: Income Residential Area $10,000 $19,999 $20,000 S39,999 $40,000 and over Urban Suburban Farm 10: 140 30 6) 15 i0 210 If a single buyer is selected from among all buyers in the region in which the survey is conducted, what is the probability that the buyer (report your answer to 4 decimal places, using conventional rounding rules) is from the urban area? has an income or less than $40,000 per year? has an income between $20,000 $39,999 per year, given that the buyer lives in a suburban area? lives in an urban area, given that the buyer has an income of at least $40,000 per year (report your answer to 4 declmal places, using conventional rounding rules) (report your answer to 4 decimal places, using conventional rounding rules) (report your answer to 4 decimal places, using conventional rounding rules) .is from a farming area or has an income of $10,000 $19,999 per year? (report your answer to 4 decimal places, using conventional rounding rulcs) does not have an income between $20,000 $39,999 per year? (report your answer to 4 decimal places, using conventional rounding rules) . is Trom a suburban area and has an income above $39.999 per year? (report your answer to 4 decimal places, using conventional rounding rules)
Answer #1
gender 10000-19999 20000-39999 40000-over total
urban 35 105 60 200
suburban 15 140 85 240
farm 10 30 20 60
total 60 275 165 500

a)P(urban area)=200/500 =0.4000

b) P(income<40000)=(60+275)/500=0.6700

c)P(income between 20000-39999 given in suburban)=140/240=0.5833

d)P(lives in urban area given income <40000 )=(35+105)/(60+275)=0.4179

e)P(farming area o income 10000-19999)=(60+60-10)/500=0.2200

f)P(nt have income between 20000-39999)=1-275/500=0.4500

g)Probability =85/500=0.1700